\begin{thebibliography}{99}
	
\bibitem{A2005} R. Askey, Evaluation of Sylvester type determinants using
orthogonal polynomials, In: H.G.W. Begehr et al. (eds.): \emph{Advances in analysis}. Hackensack, NJ, World Scientific 2005, pp. 1-16.
	
\bibitem{BR2007} T. Boros, P. R\'ozsa, An explicit formula for singular
values of the Sylvester-Kac matrix, Linear Algebra Appl. 421 (2007), 407-416.
	
\bibitem{C2010} W. Chu, Fibonacci polynomials and Sylvester determinant of
tridiagonal matrix, Appl. Math. Comput. 216 (2010), 1018-1023.
	
\bibitem{CW2008} W. Chu, X. Wang, Eigenvectors of tridiagonal matrices of
Sylvester type, Calcolo 45 (2008), 217-233.
	
\bibitem{C1959} P.A. Clement, A class of triple-diagonal matrices for test
purposes, SIAM Rev. 1 (1959), 50-52.
	
\bibitem{EK1994} A. Edelman, E. Kostlan, The road from Kac's matrix to Kac's
random polynomials, in: J. Lewis (Ed.), Proc. of the Fifth SIAM Conf. on
Applied Linear Algebra, SIAM, Philadelpia, 1994, pp. 503-507.

\bibitem{FS1965} D.K. Faddeev, I.S. Sominskii, in: J.L. Brenner (Translator), Problems in Higher Algebra, Freeman, San Francisco, 1965.
	
\bibitem{dFMMW2013} C.M. da Fonseca, D.A. Mazilu, I. Mazilu, H.T. Williams, The eigenpairs of a Sylvester-Kac type matrix associated with a simple model for one-dimensional deposition and evaporation, Appl. Math. Lett. 26 (2013), 1206-1211.
	
\bibitem{H2005} O. Holtz, Evaluation of Sylvester type determinants using
block-triangularization, In: H.G.W. Begehr et al. (eds.): \emph{Advances in analysis}. Hackensack, NJ, World Scientific 2005, pp. 395-405.

\bibitem{HZ2019} Z. Hu, P.B. Zhang, Determinants and characteristic polynomials of Lie algebras, Linear Algebra Appl. 563 (2019), 426-439.
	
\bibitem{I2002} Kh.D. Ikramov, On a remarkable property of a matrix of Mark
Kac, Math. Notes 72 (2002), 325-330.
	
\bibitem{K1947} M. Kac, Random walk and the theory of Brownian motion, Amer.
Math. Monthly 54 (1947), 369-391.
	
\bibitem{K2013} E. K\i l\i \c c, Sylvester-tridiagonal matrix with
alternating main diagonal entries and its spectra, Inter. J. Nonlinear Sci.
Num. Simulation 14 (2013), 261-266.
	
\bibitem{KA2016} E. K\i l\i \c c, T. Ar\i kan, Evaluation of spectrum of $2$-periodic tridiagonal-Sylvester matrix, Turk. J. Math. 40 (2016), 80-89.
	
\bibitem{M1923} T. Muir, \emph{The Theory of Determinants in the Historical
Order of Development}, vol. II, Dover Publications Inc., New York, $1960$
(reprinted)
	
\bibitem{OJ2017} R. Oste, J. Van den Jeugt, Tridiagonal test matrices for
eigenvalue computations: Two-parameter extensions of the Clement matrix, J.
Comput. Appl. Math. 314 (2017), 30-39.
	
\bibitem{R1957} P. R\'ozsa, Remarks on the spectral decomposition of a
stochastic matrix, Magyar Tud. Akad. Mat. Fiz. Oszt. K\"ozl. 7 (1957),
199-206.
	
\bibitem{S1926} E. Schr\"odinger, Quantisierung als Eigenwertproblem III,
Ann. Phys. 80 (1926), 437-490.
	
\bibitem{S1854} J.J. Sylvester, Th\'eor\`eme sur les d\'eterminants de M.
Sylvester, Nouvelles Ann. Math. 13 (1854), 305.
	
\bibitem{TT1991} O. Taussky, J. Todd, Another look at a matrix of Mark Kac,
Linear Algebra Appl. 150 (1991), 341-360.
	
\bibitem{V1964} I. Vincze, \"Uber das Ehrenfestsche Modell der
W\"arme\"ubertragung, Archi. Math XV (1964), 394-400.

\end{thebibliography}

\documentclass[12pt,a4paper]{amsart}
\usepackage{amssymb}
\usepackage{multirow}
\usepackage{graphicx}
\usepackage{amsmath,amsbsy,amsfonts,amsthm,latexsym,amsopn,amstext,amsxtra,euscript, amscd,mathrsfs}
\usepackage[noadjust]{cite}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}[section]

\newtheorem{conjecture}{Conjecture}

\theoremstyle{definition}

\theoremstyle{remark}

\numberwithin{equation}{section}

\numberwithin{table}{section}

\numberwithin{figure}{section}

\begin{document}

\title[An observation on the determinant of a Sylvester-Kac type matrix]{An observation on the determinant of a Sylvester-Kac type matrix}

\author{Carlos M. da Fonseca}
\address{Kuwait College of Science and Technology, Doha District, Block 4,
		P.O. Box 27235, Safat 13133, Kuwait}
\email{c.dafonseca@kcst.edu.kw}
\address{University of Primorska, FAMNIT, Glagoljsa\v ska 8, 6000 Koper, Slovenia}
\email{carlos.dafonseca@famnit.upr.si}
\author{Emrah K\i l\i \c{c}}
\address{TOBB University of Economics and Technology, Mathematics
		Department, 06560 Ankara, Turkey}
\email{ekilic@etu.edu.tr}
\subjclass[2000]{15A18, 15A15}
\date{\today }
\keywords{Sylvester-Kac matrix, Clement matrix, determinant, eigenvalues}
\begin{abstract}
Based on a less-known result, we prove a recent conjecture concerning the determinant of a certain Sylvester-Kac type matrix and consider an extension of it. 
\end{abstract}

\maketitle

\section{The Conjecture}

Quite recently, in order to find formulas for the determinants of some Lie algebras,  Z. Hu and P.B. Zhang proposed in \cite{HZ2019} the following conjecture.
\begin{conjecture} \label{conj}
The determinant of the matrix $(n+1)\times (n+1)$ 
\begin{equation*}\label{SK1}
J_{n}(z_0,z_1)=\left(
	\begin{array}{cccccc}
	z_0+nz_1 & 1 &  &  &  &  \\
	n & z_0+(n-2)z_1 & 2 &  &  &  \\
	& n-1 & z_0+(n-4)z_1 & \ddots  &  &  \\
	&  & \ddots  & \ddots  & n-1 &  \\
	&  &  & 2 & z_0-(n-2)z_1 & n \\
	&  &  &  & 1 & z_0-nz_1 
	\end{array}
	\right) \,
\end{equation*}
is 
$$
\prod_{k=0}^{n}\left( z_0-(n-2k)\sqrt{z_1^2+1}\right)\, .
$$
\end{conjecture} 

Notice that Conjecture \ref{conj} is equivalent to state that the eigenvalues of  $J_{n}(0,z_1)$ are 
$$
\pm (n-2k)\sqrt{z_1^2+1}\, , \quad \mbox{for $k=0,1,\ldots,\lfloor n/2\rfloor$.}
$$

The matrix $J_{n}(z_0,z_1)$ can be easily identified as an extension of the so-called Sylvester-Kac matrix. In fact, setting $z_1=0$ we find the characteristic matrix of the  Sylvester-Kac matrix, also known as Clement matrix, 
$$\left(
\begin{array}{ccccc}
0 & 1 &&& \\
n & 0 & 2 && \\
& n-1 & 0 & \ddots & \\
&& \ddots & \ddots & n \\
&&& 1 & 0 \\
\end{array}
\right)\, .
$$ 
The characteristic polynomial of this  matrix (that is, $\det J_{n}(x,0)$) was first conjecture in \cite{S1854}, by the $19$th century British mathematician James Joseph Sylvester celebrated,  among other facets, as the founder of the American Journal of Mathematics, in $1878$.

A fully comprehensive list of results on the different proofs for Sylvester's conjecture and the eigenpairs of non-trivial extensions of the Sylvester-Kac matrix  can be found in \cite{A2005,BR2007,C2010,CW2008,C1959,EK1994,dFMMW2013,FS1965, H2005,I2002, K1947,K2013,KA2016,M1923,OJ2017,R1957,S1926,TT1991,V1964}. 

The aim of this short note is to prove Conjecture \ref{conj} based on a result by W. Chu in \cite{C2010}. We also provide a general result containing other particular known determinants.

\section{An extension to Sylvester-Kac matrix}

In $2010$, cleverly based  on two generalized Fibonacci sequences, W. Chu proved the following theorem.

\begin{theorem}[\cite{C2010}] \label{TheoremGeneral}
	The determinant of the matrix $(n+1)\times (n+1)$ 
	\begin{equation*}\label{SK1}
	M_{n}(x,y,u,v)=\left(
	\begin{array}{cccccc}
	x & u &  &  &  &  \\
	nv & x+y & 2u &  &  &  \\
	& (n-1)v & x+2y & \ddots  &  &  \\
	&  & \ddots  & \ddots  & n-1 &  \\
	&  &  & 2v & x-(n-1)y & nu \\
	&  &  &  & v & x+ny 
	\end{array}
	\right) \,
	\end{equation*}
	is 
	$$
	\prod_{k=0}^{n}\left( x+\frac{ny}{2}+\frac{n-2k}{2}\sqrt{y^2+4uv}\right)\, .
	$$
\end{theorem}  

Of course, the formula for the determinant in Theorem \ref{TheoremGeneral} can be rewritten as
$$
\prod_{k=0}^{\lfloor n/2\rfloor}\left(\left( x+\frac{ny}{2}\right)^2-\frac{(n-2k)^2}{4}(y^2+4uv)\right)\, .
$$

Now setting $x=z_0+nz_1$, $y=-2z_1$, and $u=v=1$, we prove immediately Conjecture \ref{conj}. 

Moreover, in the spirit of \cite{A2005,dFMMW2013,H2005}, using Theorem \ref{TheoremGeneral}, we can also conclude the following theorem.

\begin{theorem}
The eigenvalues of  
	\begin{equation*}
	M^\pm_n(a,b,r) = \left( \begin{array}{ccccccc}
	n a r & b &&&&&\\
	na & ((n-1)a \pm b)r & 2b &&&& \\
	& (n-1)a & ((n-2)a\pm 2b)r &  3b &&& \\
	&& (n-2)a & \ddots & \ddots  && \\
	&&& \ddots & \ddots & nb &\\
	&&&& a & \pm nb r & \\
	\end{array} \right) \label{M} .
	\end{equation*}
are
$$\frac{1}{2} \left(n r(a\pm b)+(n-2k)\sqrt{4ab+r^2(a\mp b)^2}\right)\, ,$$ for $k=0,1,\ldots,n$....
\end{theorem}

\end{document}